OpenStudy (anonymous):

T - (-9) = 4 T=13 is that right?? If not please explain

OpenStudy (anonymous):

yes. 13 is the correct answer

jimthompson5910 (jim_thompson5910):

T - (-9) = 4 is the same as T + 9 = 4

OpenStudy (anonymous):

if u plug 13 into T the 2 negatives cancel and become postive

jimthompson5910 (jim_thompson5910):

since subtracting a negative is the same as adding

OpenStudy (anonymous):

Alright thx

OpenStudy (anonymous):

welcome

jimthompson5910 (jim_thompson5910):

Let's check if 13 is the answer T + 9 = 4 13 + 9 = 4 ... replace T with 13 22 = 4 ... this equation is false, so T = 13 is not the answer

OpenStudy (anonymous):

Ohh they just told me it was correct

OpenStudy (anonymous):

Lol. Ok. Can you explain it to me

jimthompson5910 (jim_thompson5910):

well that check says otherwise

jimthompson5910 (jim_thompson5910):

How would you isolate T in T + 9 = 4

OpenStudy (anonymous):

T=-4

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

what must you do to both sides of T + 9 = 4 to get T all by itself?

jimthompson5910 (jim_thompson5910):

hint: you want to move that +9 over

OpenStudy (anonymous):

Plus 9 to both sides

jimthompson5910 (jim_thompson5910):